OpenStudy (anonymous):

Solve the inequalities (x^2-x-6)/(x^2+4x-5)>=0.

OpenStudy (anonymous):

x>=3 -2=<x<1 x<-5

OpenStudy (anonymous):

Now What you should do here is obtain factors of both numerator and denominator

OpenStudy (anonymous):

\[\frac{x^2 -x -6 }{x^2 + 4x -5} \ge 0\]

OpenStudy (anonymous):

\[\frac{(x -2)(x +3)}{(x + 5)(x - 1)} \ge 0\]

OpenStudy (anonymous):

then plot them on number line ... and determine the sign of the whole function now this the case of > 0 ... so you should only include those intervals which give positive value of functions

OpenStudy (anonymous):

hmm how do i see whether it has to include infinity or not?

OpenStudy (anonymous):

sorry typo its (x + 2)(x - 3) in numerator

OpenStudy (anonymous):

Choice given is in the attachment.

myininaya (myininaya):

\[\frac{x^2-x-6}{x^2+4x-5}=\frac{(x-3)(x+2)}{(x+5)(x-1)} \ge 0\] so the areas we test around will be -5,-2,1, and 3 ----|-----|----|------|----- -5 -2 1 3 ands its =0 when x=3 and x=-2 (so let's include that in our solution) let \[f(x)=\frac{(x-3)(x+2)}{(x+5)(x-1)}\] choose a number before -5 like -6 plug into f(x) f(-6)=(-)(-)/(-)(-)=+ so (-inf,-5) is a part of out solution now lets choose a number between -5 and -2 like -4 f(-4)=(-)(-)/(+)(-)=- so (-5,-2) is not a part of our solution now choose number between -2 and 1 like 0 f(0)=(-)(+)/(+)(-)=+ sp (-2,1) is a part of our solution now choose a number between 1 and 3 like 2 f(2)=(-)(+)/(+)(+)=- so (1,3) is not a part of our solution now last interval to check (3,inf) choose a number in that interval like 4 f(4)=(+)(+)/(+)(+) so (3,inf) is a part of our solution

myininaya (myininaya):

so the answer is (-inf,-5) U [-2,1) U [3,inf) do you remember why i include -2 and 3? because it saids =0 also

OpenStudy (anonymous):

thanks a lot!!